OpenStudy (anonymous):

In a classroom, the students are 10 boys and 18 girls. If one student is selected at random, find the probability that the student is a boy.

jimthompson5910 (jim_thompson5910):

P(Boy) = (# of boys)/(# of students total) = 10/(10+18) = 10/28 = 5/14 So the probability is 5/14